OpenStudy (anonymous):

How would you plug this function: 3(x square root) , whose upper limit is 10 and lower one is 0, and this one: ^3(x "square root sign) - 1/4(^3(x "square root sign")) into the arc length formula? There derivatives are originially from 2(x^3/2), and 3/4(x^4/3) - 3/8(x^2/3)+7.. @Calculus1

OpenStudy (turingtest):

can you use the equation feature please? Trying to read this is hurting my eyes.

OpenStudy (saifoo.khan):

\[Turing-Test.\]

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=J6ZWlDks0nQ

OpenStudy (turingtest):

SAIFOO-SAN-KHAN

OpenStudy (saifoo.khan):

San?? o_O from where did that came from?

OpenStudy (turingtest):

like master saifoo khan san turing test san you know...

OpenStudy (turingtest):

like sensei, you know

OpenStudy (saifoo.khan):

Ohh!! now i got it.. ^_^ thanks. =)

OpenStudy (turingtest):

is this the music you do math to? pretty mellow vibe

OpenStudy (saifoo.khan):

Nope, i go for random music. currently this once's on top! =D

OpenStudy (turingtest):

oh, cuz this is what I listen to http://www.youtube.com/watch?v=OFSrINhfNsQ

OpenStudy (saifoo.khan):

LOL.

OpenStudy (turingtest):

:)

OpenStudy (anonymous):

alright hold on

OpenStudy (anonymous):

is that your pic Turing?

OpenStudy (saifoo.khan):

i guess so.

OpenStudy (turingtest):

yup that's me

OpenStudy (turingtest):

that's you saifoo?

OpenStudy (saifoo.khan):

Yes, me and my camera! =D

OpenStudy (anonymous):

Nice hairs! :D

OpenStudy (saifoo.khan):

Agree^

OpenStudy (turingtest):

Heh, I was at a hippie commune in that pic. Only hippie physics major there ;) what ethinicity/nationality are you?

OpenStudy (anonymous):

\[y=2x ^{3/2}\] is the given curve where 0 =a and 10=b and \[y = 3/4x ^{4/3}-3/8x ^{2/3}+7\] where 1 is a and 8 is b.

OpenStudy (saifoo.khan):

Pakistani. you?

OpenStudy (anonymous):

I am an Indian .. and you ?

OpenStudy (turingtest):

All American ,but I live in mexico

OpenStudy (turingtest):

ok let's do our buddy's problem, now that it's posted correctly

OpenStudy (anonymous):

I'm trying to find the arc lengths of both functions, I already found the derivatives, I'm just unsure of how to plug that back into the function.

OpenStudy (anonymous):

OH mostly Mexican want to leave in America.. any felony records in states ? :P

OpenStudy (anonymous):

ok thanks.

OpenStudy (anonymous):

what is a and b ?

OpenStudy (anonymous):

integration ?

OpenStudy (anonymous):

a stand for the lower limit and b stands for the upper limit.

OpenStudy (anonymous):

^^Yes

OpenStudy (anonymous):

so this are two separate problems of finding the area?

OpenStudy (anonymous):

*these

OpenStudy (turingtest):

these are after you've taken the derivative, correct?

OpenStudy (anonymous):

yes, that is correct.

OpenStudy (anonymous):

\[ \int_0^{10} 2x ^{3/2} dx \]

OpenStudy (turingtest):

so why do you have y= and not dy/dx= ???

OpenStudy (anonymous):

No actually I was wrong, both functions are before doing the derivative

OpenStudy (turingtest):

ok let's take the derivative then...

OpenStudy (anonymous):

whatever you are just flustered yourself.....

OpenStudy (turingtest):

\[dy/dx=3x^{1/2}\]\[dy/dx=x^{1/3}-{x^{-1/3}\over4}\]are the derivatives then. do we agree?

OpenStudy (turingtest):

arc length formula:\[\int\limits_{a}^{b} \sqrt{1+(dy/dx)^2}\]for the first prob\[\int\limits\limits_{0}^{10} \sqrt{1+(3x^{1/2})^2}dx=\int\limits\limits_{0}^{10} \sqrt{1+9x}dx\]can you integrate that yourself?

OpenStudy (anonymous):

^ Yes that is what I got, except you have to bring each of the exponents down and square them. For your other question, I got \[\sqrt{91}\]

OpenStudy (turingtest):

the above integral\[\int\limits_{0}^{10}(1+9x)dx={2\over27}(1+9x)^{3/2}\]from 0 to 10 is\[{2\over27}(91^{3/2}-1)\]

OpenStudy (turingtest):

sorry that should be (1+9x)^(1/2) in the integral.

OpenStudy (anonymous):

Yes this is what I got and the decimal approximation to it would be 64.228

OpenStudy (anonymous):

Okay, now I think I can figure out the second one. I'll let you know if I have any issues. Thanks alot.

OpenStudy (turingtest):

the other becomes\[dy/dx=x^{1/3}-{1\over4}x^{-1/3}\]\[(dy/dx)^2=x^{2/3}+x^{-2/3}-1/2\]so arc length becomes\[\int\limits_{1}^{8}\sqrt{x^{2/3}+x^{-2/3}+{1\over2}}dx=\int\limits_{1}^{8}\sqrt{x^{2/3}+x^{-2/3}+{1\over2}}dx\]or at least that's what I got, but I can't seem to integrate it, so maybe you can help me with this one when you figure it out

OpenStudy (anonymous):

alright, I'll look at it later and get back to you.

OpenStudy (turingtest):

actually (dy/dx)^2 should be \[x^{2/3}+{1\over16}x^{-2/3}-{1\over2}\]my mistake

OpenStudy (turingtest):

so\[\int\limits_{1}^{8}\sqrt{x^{2/3}+{1\over16}x^{-2/3}+{1\over2}}dx\]should be the formula...

OpenStudy (turingtest):

I want to factor what's in the radical into (a+b)^2 but that doesn't seem possible...

OpenStudy (turingtest):

I think I got it...

OpenStudy (turingtest):

\[\int\limits_{1}^{8} \sqrt{x^{2/3}+{1\over2}+{1\over16}x^{-2/3}}dx=\int\limits_{1}^{8} \sqrt{x^{-2/3}(x^{4/3}+{1\over2}x^{2/3}+{1\over16})}dx\]\[=\int\limits_{1}^{8}\sqrt{x^{-2/3}(x^{2/3}+{1\over4})^2}dx=\int\limits_{1}^{8}x^{-1/3}(x^{2/3}+{1\over4})dx\]\[=\int\limits\limits_{1}^{8}x^{1/3}+{1\over4}x^{-1/3}dx={3\over4}x^{4/3}+{3\over8}x^{2/3}\]I'm prettty sure you can evaluate that from 1 to 8 yourself, right?

OpenStudy (turingtest):

I got 93/8